OpenStudy (anonymous):

Why is pi(2r)^2(2h) the same as (8)pi(r)^2(h)? It just doesn't make sense to me.. :/

OpenStudy (anonymous):

\[(2r)^2=2r\times 2r=4r^2\] \[4r^2\times 2h=8r^2h\]

OpenStudy (anonymous):

But I thought you couldn't multiply different variables?

OpenStudy (p0sitr0n):

you can, you just cant add their degrees together

OpenStudy (anonymous):

Ohhh, thank you

OpenStudy (anonymous):

you can multiply anything you like

OpenStudy (anonymous):

when you multiply numbers together you actually do a multiplication, like \(4\times 2=8\) when you multiply variables together, you just write then next to each other like \(r^2\times h=r^2h\)

OpenStudy (anonymous):

So you can always multiply coefficients as long as you then at the end write the variables next to each other?